# ShareChat: A Dataset of Chatbot Conversations in the Wild

### v1

@@ Appendix A Appendix
A.1 Turn length Distribution
@@
A.2 Toxicity Analysis Results
Leveraging the multi-label capabilities of the Detoxify model trained on the Jigsaw Toxic Comment Classification dataset, we analyzed six distinct dimensions of harmful content: toxicity, which is a broad category for rude or disrespectful comments, severe toxicity, obscenity, threat, insult, and identity attack. Table 4 presents the mean probability scores for each of these six dimensions by platform. The final column, “Ratio Toxic,” reports the percentage of turns where the general toxicity score exceeded the standard threshold of 0.1.
@@ A.2 Toxicity Analysis Results
| (%) | Detox | OpenAI | Detox | OpenAI | Detox | OpenAI | Detox | OpenAI | Detox | OpenAI | Detox | OpenAI |
@@
A.3 Detailed Analysis of User Request Topic Distribution
@@
A.4 Conversation Completeness Analysis
We provided the prompts used for the conversation completeness analysis with Qwen3-8B, adapted from the DeepEval framework in this section. To handle conversations that exceed the 40,960-token context limit of Qwen3-8B, we apply token-based truncation to assistant responses while preserving all user content, and we exclude conversations requiring severe truncation (retaining less than 80% of content) to maintain evaluation quality.
@@ A.4 Conversation Completeness Analysis
A.5 Additional Temporal Analysis: Text Length and User Response Time

### v2

@@ Appendix A Appendix
A.1 Data collection
Table 3 details the target URLs used for extraction and the data collection timeframe for each platform.
| Platform | Share URL Format | Collection Period |
| ChatGPT | chatgpt.com/share/* | May 2023 to Aug 2025 |
| Perplexity | perplexity.ai/search/* | Apr 2023 to Oct 2025 |
| Grok | grok.com/share/* | Dec 2024 to Oct 2025 |
| Gemini | gemini.google.com/share/* | Apr 2024 to Sep 2025 |
| Claude | claude.ai/share/* | Not available |
A.2 Platform-Specific Interaction and Metadata Attributes
Different platforms also reflect distinct design philosophies, ChatGPT and Gemini offer general-purpose conversational assistance. Perplexity focuses on search-oriented information retrieval. Grok targets social media-integrated question answering. Claude specializes in technical and reasoning support. These structural differences align with intended representative analyses and shape the types of user interactions preserved in ShareChat.
As shown in Table 2, while ChatGPT and Gemini follow standard conversational formats, other platforms expose specialized content structures. Claude is unique in preserving “thinking blocks” and versioned code artifacts, enabling the analysis of intermediate reasoning and iterative coding processes. In contrast, Perplexity and Grok structure responses around retrieval. Perplexity organizes content into answers, sources, and images with inline citations, while Grok features a dedicated pane for web and social media sources (X posts). These structural differences allow for distinct lines of inquiry: technical workflow analysis for Claude versus citation accuracy and source integration studies for Perplexity and Grok.
A.3 Turn length Distribution
@@
A.4 Toxicity Analysis Results
Leveraging the multi-label capabilities of the Detoxify model trained on the Jigsaw Toxic Comment Classification dataset, we analyzed six distinct dimensions of harmful content: toxicity, which is a broad category for rude or disrespectful comments, severe toxicity, obscenity, threat, insult, and identity attack. After filtering conversations supported by the model, resulting in a filtered dataset of 104,107 (72.9% of the original corpus) across all five platforms. Table 4 presents the mean probability scores for each of these six dimensions by platform. The final column, “Ratio Toxic,” reports the percentage of turns where the general toxicity score exceeded the standard threshold of 0.1.
@@ A.4 Toxicity Analysis Results
| (%) | Detox | Open | Detox | Open | Detox | Open | Detox | Open | Detox | Open | Detox | Open |
@@
A.5 Detailed Analysis of User Request Topic Distribution
@@
A.6 Conversation Completeness Analysis
The pipeline implements a three-stage workflow adapted from the DeepEval444https://github.com/confident-ai/deepeval?tab=readme-ov-file conversation completeness metric framework. First, we extract user intentions from each conversation by prompting Qwen3-8B yang2025qwen3 with a temperature of 0.7 to identify the distinct goals or information needs expressed in the user turns, yielding a chronologically ordered list of intentions per conversation. Second, for each identified intention, we construct an evaluation prompt that presents the full conversation context alongside the specific intention and instruct Qwen3-8B to classify whether that intention was satisfied; the model produces a categorical verdict of complete (the assistant fully addressed the request), partial (the assistant provided some relevant information but the goal was incompletely met), or incomplete (the assistant failed to address the request or the conversation ended before meaningful progress). Finally, we aggregate these intention-level verdicts into a conversation-level completeness score by computing the weighted proportion of intentions, where complete verdicts contribute 1.0, partial verdicts contribute 0.5, and incomplete verdicts contribute 0.0. This three-way categorical framework allows us to capture nuanced degrees of conversational success beyond binary outcomes. We apply this pipeline uniformly across all five platforms in ShareChat, enabling direct cross-platform comparisons of how effectively different assistants satisfy user needs over the course of extended interactions. During manual coding, we observe that mis-classification often occurred when users pasted text, such as articles or messages, without explicit instructions or questions. In these instances, the LLM generate responses based solely on an inferred understanding of the content.
Figure 10 shows the conversation-level completeness scores, where ChatGPT, Claude, Grok, and Perplexity all achieve median scores of 1.0, while Gemini’s median is 0.833. Figure 11 shows the number of interactions across platforms.
The prompts used are as follows. To handle conversations that exceed the 40,960-token context limit of Qwen3-8B, we apply token-based truncation to assistant responses while preserving all user content, and we exclude conversations requiring severe truncation (retaining less than 80% of content) to maintain evaluation quality.
@@ A.6 Conversation Completeness Analysis
A.7 Additional Source Grounding Analysis
To understand the retrieval intensity of search-enabled platforms, we analyze the distribution of source counts per conversation. Figure 12 shows that Grok typically uses very few sources per conversation, whereas Perplexity exhibits a long-tailed distribution in which many conversations draw on dozens of sources.
A.8 Additional Temporal Analysis
We expand on the temporal dynamics of the dataset by examining both the diurnal rhythms of user activity and the granular relationship between content length and reading time. First, to understand when users are most active, Figure 13 presents the normalized hourly activity for each platform, with timestamps adjusted to the users’ local time. Second, we investigate whether longer model outputs consistently drive longer user engagement. Figure 14 plots the raw distribution of user response times against LLM response length. The scatter plot reveals a high degree of variance, with Pearson correlations near zero for both platforms (ChatGPT: ; Grok: ). This indicates that while aggregate trends exist, model verbosity is a poor predictor of dwell time.